Project

General

Profile

NDOS Commissioning Meeting 2011/04/21

Attendance: Jon, Gavin, Enrique, Rafael, Susan, Mat, Mark Mathis, Andrew, Gary, Jarek, Jan, Nick, Zeb, Leon, Peter, Mark Messier

Agenda:

  1. NDOS update
  2. DAQ/Electronics Update & Plans
  3. DCS Status & Plans
  4. DataCheck update
  5. DSO Studies
  6. AOB

NDOS update

  • Mat:
    • top-off has been going on all week; as of yesterday, the middle and bottom modules of the MC and 3rd di-blocks have been finished; next week we will focus on the top row
    • FEBs have been installed on MC; need to be checked out (will ask Brian to do this while on shift)
    • FEBs have been installed on 5 planes of 3rd-diblock (~1/3 of what is planned); another 6 or 7 will come from Harvard, the rest from the test-stand when APDs are available
    • APDs to be installed on MC tomorrow (horizontals)
    • next week we will begin recovery work for APDs in 1st two di-blocks
    • there are ~30 cable-ectomy candidates; Leon is looking for extra cable-ribbon supplies at CalTech
    • CAP board installation has been going on this week. All horizontal APDs on top row and all APDs on 1st diblock are done. This work will continue tomorrow and then resume next Tuesday.
    • help is still needed next Wed.-Fri. for various tasks out at the NDSB

DAQ/Electronics Status & Plans

  • Andrew:
    • are we globally synchronized in time? Has the NuMI peak in the timing moved?
    • we have been running in the new TDU master-slave configuration since Monday without problems.
    • would like to get multiple timing chains introduced (can slave drive two outputs?)
    • next (TDU): move master to MINOS building

DCS Status & Plans

  • Chad: presently installing a patch panel in rack two and replacing the immersion probe RTDs used in the rack with my newly soldered variety
  • Andrew: CD has full EPICS system running on NOvA DAQ test-stand; they are able to read out temperatures, and have a GUI to show this data. Hoping to have something functional and ready-to-go by sometime next week

DataCheck Update

DSO Scan Update

AOB

  • Peter: anyone figure out what was wrong with dcm-2-3, feb 1? This gave a certes lock error during hardware config, so it was eventually removed from hardware list (DSO scan produced a mask for this FEB). Needs to be further investigated.